The surface area of a sphere inscribed in a regular tetrahedron is 144pi centimeter squared.
1. What is the radius of the sphere?

2. What is the altitude of the tetrahedron?

3. If a sphere is inscribed in a cube of side 12cm, what is the volume of the sphere?
1
Expert's answer
2020-12-16T18:53:44-0500

1 . The surface area of a sphere of radius "r" is

"S=4""\pi r^2"

then,

"r=\sqrt{\frac{S}{4\pi}}"

Given "S=144\pi" "cm^2"

"r=\sqrt{\frac{144\pi cm^2}{4\pi}} = \frac{12}{2} = 6" "cm"

2 . For a regular tetrahedron of edge length "a" :

Altitude "h=\sqrt{\frac{2}{3}}a"

"a=2r\sqrt6" "=12\sqrt6"

"h=\sqrt\frac{2}{3}\times12\sqrt6 = 24"

3 . The diameter of the sphere would be equal to the side of the square .

"d=12cm"

then,

"r=6cm"

So,

"V=\frac{4}{3} \times \pi r^3"

"V=\frac{4}{3} \times \pi \times216"

"V=904.32"